OpenStudy (anonymous):

If w = <-8, 15>, find ||w||. 7 17 23

OpenStudy (anonymous):

Sqrt(7)

OpenStudy (anonymous):

sqrt(23)

OpenStudy (johnweldon1993):

\[\large ||w|| = \sqrt{w_x^2 + w_y^2}\]

OpenStudy (anonymous):

I got 23

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

\[\large ||w|| = \sqrt{(-8)^2 + 15^2} = \sqrt{64 + 225} = \sqrt{289} = ?\]

OpenStudy (anonymous):

17

OpenStudy (johnweldon1993):

There we go!